# Improving Nonpreemptive Multiserver Job Scheduling with Quickswap

### v1

@@ Abstract
Modern data center workloads are composed of multiserver jobs, computational jobs that require multiple CPU cores in order to run. A data center server can run many multiserver jobs in parallel, as long as it has sufficient resources to meet their individual demands. However, multiserver jobs are generally stateful, meaning that job preemptions incur significant overhead from saving and reloading the state associated with running jobs. Hence, most systems try to avoid these costly job preemptions altogether. Given these constraints, a scheduling policy must determine what set of jobs to run in parallel at each moment in time to minimize the mean response time across a stream of arriving jobs. Unfortunately, simple non-preemptive policies such as First-Come First-Served (FCFS) may leave many cores idle, resulting in high mean response times or even system instability. Our goal is to design and analyze non-preemptive scheduling policies for multiserver jobs that maintain high system utilization to achieve low mean response time.
One well-known non-preemptive scheduling policy, Most Servers First (MSF), prioritizes jobs with higher core requirements and is known for achieving high resource utilization. However, MSF causes extreme variability in job waiting times, and can perform significantly worse than FCFS in practice. To address this issue, we propose and analyze a class of scheduling policies called Most Servers First with Quick Swap (MSFQ) that performs well in a wide variety of cases. MSFQ reduces the variability of job waiting times by periodically granting priority to other jobs in the system. We provide both stability results and an analysis of mean response time under MSFQ to prove that our policy dramatically outperforms MSF in the case where jobs either request one core or all the cores. In more complex cases, we evaluate MSFQ in simulation. We show that, with some additional optimization, variants of the MSFQ policy can greatly outperform MSF and FCFS on real-world multiserver job workloads.
@@ 1 Introduction
Modern data centers serve multiserver jobs that occupy multiple servers or CPU cores simultaneously [26, 35, 14, 13]. Each multiserver job has an associated server need, the number of servers the job requires to run, and job size, the amount of time the job must run to be completed. A set of multiserver jobs can run in parallel, but only if the system has enough dedicated servers for each job. A data center scheduling policy must select which jobs to run in parallel at every moment in time. Given a fixed number of servers, $k$, a common goal is to design a scheduling policy that minimizes the mean response time across jobs in a stream of arriving multiserver jobs — the average time from when a job arrives to the system until it is completed.
There are two central difficulties in designing performant scheduling policies for multiresource jobs. First, modern datacenter workloads generally exhibit large variability in both the server needs and sizes of their jobs [35]. As a result, it is usually impossible to utilize all available servers using the set of multiserver jobs currently in the system. In general, leaving more servers unutilized on average will lead to higher mean response time or even system instability. Unfortunately, maximizing the number of utilized servers at a specific moment in time requires solving a knapsack problem instance. It is even more difficult, then, to maximize the utilization of the available servers as jobs enter and exit the system over time.
Second, modern multiserver jobs are generally stateful, meaning that job preemptions require persisting and/or reloading a significant amount of program state [31]. As a result, job preemptions or migrations can take a prohibitively long amount of time to perform. Due to this overhead, data centers typically employ non-preemptive scheduling policies that avoid costly preemptions altogether. Given these two constraints, this paper designs and analyzes new, non-preemptive scheduling policies that aim to minimize the mean response time across a stream of multiserver jobs.
@@ 1.1 Prior Approaches
Much of the prior work on multiserver job scheduling uses frequent job preemptions to ensure that server utilization remains high as jobs enter and leave the system [11, 24, 20, 19]. These preemptive policies are of limited utility when processing the stateful jobs that are common in data centers.
@@
First-Come First-Served (FCFS) is a naïve non-preemptive policy that serves jobs in arrival order until the system runs out of available servers. For example, when a job with a large server need reaches the front of the queue, the system may not have enough available servers to fit this job in service. FCFS stops scheduling additional jobs at this point, even if other jobs in the queue could fit into service. This phenomenon, known as Head-of-the-Line blocking, causes FCFS to underutilize servers, resulting in high mean response time. Although FCFS is a simple policy, analyzing it has been proven difficult due to its dependence on the random arrival order of jobs with different server needs. Only recently, [23] derived mean response time bounds that confirm the empirical observation that FCFS performs poorly in practice.
@@
Most Servers First (MSF) [34, 5, 20] is a non-preemptive policy that prioritizes jobs with larger server needs. Specifically, whenever the system has available servers, jobs are considered in descending server need order. Jobs that find their required number of servers being put in service successively. To understand both the benefits and the pitfalls of MSF, consider an example where jobs either need either 1 server or $k$ servers. We refer to this case as the one-or-all case for multiserver jobs. In this case, MSF serves jobs in two alternating phases. First, MSF serves $k$-server jobs until none remain in the system. Then MSF serves 1-server jobs until none remain before returning to serve $k$-server jobs. We will show in Section 5.1 that, by switching between these two phases, MSF achieves optimal long-run average server utilization in the one-or-all case.
@@
While one might hope that MSF leverages its high server utilization to achieve low mean response time, we also find that MSF takes an increasingly long time to switch phases as the job arrival rate increases (see Section 5.3). This creates a feedback loop in the system whereby many 1-server jobs accumulate while the system processes $k$-server jobs, leading to a long period of serving 1-server jobs during which many $k$-server jobs will accumulate. The two things to note about this process are that (i) despite its name, MSF can spend long periods of time giving priority to 1-server jobs over $k$-server jobs and (ii) because the class of jobs not in service accumulates quickly, there are almost always a large number of jobs in the system under MSF. Figure 1 illustrates this problem via simulations that track the number of jobs in the system under MSF for the one-or-all case. As MSF alternates between phases, jobs of the opposite class accumulate quickly in the queue. While all jobs are eventually served, this behavior ensures that a significant fraction of arriving jobs have long queueing times, leading to a high overall mean response time.
(a) Number of small jobs vs time
(b) Number of large jobs vs time
@@
Figure 1: Number of jobs in the system under MSFQ and MSFQ where there are $k$=32 servers, 90% of the job arrivals are 1-server jobs, mean job sizes are 1 for 1-server jobs and $k$-server jobs, and jobs arrive at a rate of 7.5 jobs/second.
@@
First-Fit is a variant of FCFS that avoids head-of-line blocking by continuing to examine jobs in arrival order after finding a job that does not fit in service. One might hope that this policy gets a near-optimal server utilization without the harmful periodic behavior of MSF. Unfortunately, in the one-or-all case, First-Fit has the same alternating behavior as MSF, but First-Fit spends even more time serving 1-server jobs.
@@ 1.2 A New Approach: Most Servers First with Quickswap
Our central observation is that, while MSF achieves high server utilization, it does a poor job of switching which class of jobs are in service. Specifically, in the one-or-all case, MSF ties the decision to switch from serving 1-server jobs to serving $k$-server jobs to the time required to drain all 1-server jobs from the system. This time, which is essentially a partial busy period in an $M/M/k$ queue, explodes as the arrival rate increases or as $k$ becomes large [4], allowing many $k$-server jobs to accumulate in the system. To reduce mean response time, our goal is to maintain the high utilization of MSF while shortening the time the scheduling policy takes to switch between job classes.
With this motivation in mind, we propose a new class of policies called Most Servers First with Quickswap (MSFQ), which is designed to improve the performance of MSF in the one-or-all case. Unlike MSF, which tries to drain all 1-server jobs from the system before switching, MSFQ switches to serving $k$-server jobs whenever the number of 1-server jobs in the system falls below a threshold, $\ell$. That is, when the number of 1-server jobs falls below $\ell$, MSFQ stops admitting jobs into service and lets all running 1-server jobs complete. MSFQ can then begin serving $k$-server jobs. By increasing $\ell$, we can shorten the time MSFQ requires to switch phases, reducing the mean response time compared to MSF. Furthermore, we will prove that MSFQ maintains the high server utilization achieved by MSF.
Figure 1 compares the number of jobs in the system under the MSF and MSFQ policies (the yellow and blue curves, respectively). Setting a sufficiently high value of $\ell$ for MSFQ ($\ell=k-1$ in this case) greatly dampens the feedback that causes jobs to accumulate under MSF. As a result, an MSFQ policy can achieve a much lower mean response time than MSF. We will also show that MSFQ policies are much better at balancing the mean response time between different job classes. While our MSFQ policy is tailored specifically to the one-or-all case, we will also explore generalizations of this policy to an arbitrary number of job classes.
@@ 1.3 Contributions
Inspired by the high server utilization of the MSF policy, this paper formally defines and analyzes the class of MSFQ policies for scheduling multiserver jobs. Our analysis has two main goals. First, we aim to prove stability results about MSFQ policies. We say the system is stable under a scheduling policy if the policy achieves sufficiently high server utilization such that the mean number of jobs in the system and the mean response time across jobs are both finite. Second, we will analyze the mean response time under MSFQ policies and their variants both in theory and in simulation to show the advantages of these policies compared to prior approaches. Specifically, the contributions of this paper are as follows.
@@
  • In Section 4.2 we introduce the MSFQ policy, which uses the Quickswap mechanism to force the policy to switch phases faster. Crucially, in Section 5.1 we show that any MSFQ policy matches the server utilization of MSF by proving that MSFQ is throughput-optimal. Here, throughput-optimality means that MSFQ will stabilize the system whenever the system can be stabilized.
  • Then, in Section 5.3, we analyze the mean response time under MSFQ by approximating the Laplace Transform of the phase durations and number of jobs at the beginning of each phase. While the MSFQ system resembles a polling system (See Section 2.3), MSFQ cannot be analyzed using existing results from the polling literature. Hence, our analysis of MSFQ also represents a new contribution to the extensive body of work on polling systems.
  • Finally, Section 6 examines generalizations of MSFQ to real-world settings where jobs’ server needs can vary widely. We evaluate two generalizations of MSFQ, called Static Quickswap and Adaptive Quickswap, and evaluate these policies in simulation using traces from the Google Borg cluster scheduler [35]. Our simulations show that a well-designed Quickswap policy can improve mean response time by orders of magnitude. Furthermore, Quickswap policies tend to achieve an equitable mean response time between the job classes as compared to a less fair priority policy like MSF.
@@ 2.1 Systems for Multiserver Job Scheduling
Modern data centers schedule multiserver jobs across thousands of machines, supporting workloads with diverse server needs and job sizes [27, 35, 37]. None of these schedulers make formal performance guarantees about system stability or mean response time, generally relying on heuristics to make scheduling decisions.
SLURM [27] is an open-source cluster scheduler used in data centers and high-performance computing environments. It uses a combination of heuristics and a variant of FCFS scheduling called BackFilling. While this approach can improve resource utilization by running low-priority jobs opportunistically, it requires accurate predictions of job sizes to work well, and can therefore suffer from low server utilization in practice. Borg [35], Google’s internal resource management system for data centers, schedules batch jobs by placing incoming jobs in an FCFS queue. Once there is enough capacity to serve the next batch job, complex heuristics are used to assign the job to a specific set of servers. YARN [37], integrated with Hadoop, supports both FCFS and other heuristic policies aimed at optimizing fairness instead of stability or mean response time. Hence, many systems designed to schedule multiserver jobs stand to benefit from improved scheduling policies that are accompanied by formal performance guarantees.
@@ 2.2 Multiserver Job Scheduling in Theory
Prior work from the theory community on multiserver jobs has mostly focused on the stability and response time analysis of FCFS. The stability region of FCFS was studied in [33, 29, 1] in the case where all job sizes follow the same exponential distribution. Subsequently,  [22, 30] considered the case where jobs belong to one of two job classes, deriving explicit expressions for the stability region of FCFS. For many years, mean response time analysis of FCFS was restricted to systems with just two servers [10, 15]. However,  [23] recently derived explicit bounds on mean response time that are tight up to an additive constant. Matrix geometric approaches [2, 3] have also recently been used to characterize the performance of FCFS systems with two job classes under specific service time distributions. While FCFS is becoming well-understood, all of these analyses confirm that it can perform poorly in terms of both stability and mean response time.
There is comparatively little work on more complex and efficient policies that do not require job preemptions. For example, the well-studied MaxWeight policy is throughput optimal, but requires preemption and is computationally costly to implement in practice [28]. Other recent work on scheduling multiserver jobs has also been restricted to the case of preemptible jobs [21, 24]. There are two prominent examples of throughput-optimal, non-preemptive policies for scheduling multiserver jobs. First, Randomized Timers is a throughput-optimal policy based on MaxWeight that is non-preemptive [32]. Unfortunately, there is no known mean response time analysis of Randomized Timers, and the policy has been shown perform poorly in practice. Second, [12] recently analyzed a new class of non-preemptive policies called Markovian Service Rate (MSR) policies. An MSR policy precomputes a set of schedules with high server utilization, and switches between schedules according to a continuous-time Markov chain that is independent of the system state (i.e., queue lengths). The class of MSR policies is throughput-optimal and admits an analysis of mean response time. However, because MSR policies do not consider queue length when switching schedules, they waste capacity unnecessarily, resulting in high mean response time. We will show that MSFQ can significantly outperform MSR policies by considering queue length when switching schedules.

### v2

@@ Abstract
Modern data center workloads are composed of multiserver jobs, computational jobs that require multiple servers in order to run. A data center can run many multiserver jobs in parallel, as long as it has sufficient resources to meet their individual demands. Multiserver jobs are generally stateful, meaning that job preemptions incur significant overhead from saving and reloading the state associated with running jobs. Hence, most systems try to avoid these costly job preemptions altogether. Given these constraints, a scheduling policy must determine what set of jobs to run in parallel at each moment in time to minimize the mean response time across a stream of arriving jobs. Unfortunately, simple non-preemptive policies such as First-Come First-Served (FCFS) may leave many servers idle, resulting in high mean response times or even system instability. Our goal is to design and analyze non-preemptive scheduling policies for multiserver jobs that maintain high system utilization to achieve low mean response time.
One well-known non-preemptive scheduling policy, Most Servers First (MSF), prioritizes jobs with higher server needs and is known for achieving high resource utilization. However, MSF causes extreme variability in job waiting times, and can perform significantly worse than FCFS in practice. To address this issue, we propose and analyze a class of scheduling policies called Most Servers First with Quickswap (MSFQ) that performs well in a wide variety of cases. MSFQ reduces the variability of job waiting times by periodically granting priority to other jobs in the system. We provide both stability results and an analysis of mean response time under MSFQ to prove that our policy dramatically outperforms MSF in the case where jobs either request one server or all the servers. In more complex cases, we evaluate MSFQ in simulation. We show that, with some additional optimization, variants of the MSFQ policy can greatly outperform MSF and FCFS on real-world multiserver job workloads.
@@ 1 Introduction
Modern data centers serve multiserver jobs that occupy multiple servers simultaneously [mor-2022, tirmazi2020borg, delimitrou2014quasar, dean2008mapreduce]. Each multiserver job has an associated server need, the number of servers the job requires to run, and job size, the amount of time the job must run to be completed. A set of multiserver jobs can run in parallel, but only if the system has enough dedicated servers for each job. A data center scheduling policy must select which jobs to run in parallel at every moment in time. Given a fixed number of servers, $k$, our goal is to design a scheduling policy that minimizes the mean response time across jobs in a stream of arriving multiserver jobs — the average time from when a job arrives to the system until it is completed.
There are two central difficulties in designing performant scheduling policies for multiserver jobs. First, modern datacenter workloads generally exhibit large variability in both the server needs and sizes of their jobs [tirmazi2020borg]. As a result, it is usually impossible to utilize all available servers using the set of multiserver jobs currently in the system. In general, leaving more servers unutilized on average will lead to higher mean response time or even system instability. Unfortunately, maximizing the number of utilized servers at a specific moment in time requires solving a knapsack problem instance. It is even more difficult, then, to maximize the utilization of the available servers as jobs enter and exit the system over time.
Second, modern multiserver jobs are generally stateful, meaning that job preemptions require persisting and/or reloading a significant amount of program state [psychas2017non]. As a result, job preemptions or migrations can take a prohibitively long amount of time to perform. Due to this overhead, data centers typically employ non-preemptive scheduling policies that avoid costly preemptions altogether. Given these two constraints, this paper designs and analyzes new, non-preemptive scheduling policies that aim to minimize the mean response time across a stream of multiserver jobs.
@@ 1.1 Prior Approaches
Much of the prior work on multiserver job scheduling uses frequent job preemptions to ensure that resource utilization remains high as jobs enter and leave the system [chen.usenix, grosof-2022-pomacs, grosof-2023-serverfilling, georgiadis-2006]. These preemptive policies are of limited utility when processing the stateful jobs that are common in data centers.
@@
First-Come First-Served (FCFS) is a naïve non-preemptive policy that serves jobs in arrival order until the system runs out of available servers. For example, when a job with a large server need reaches the front of the queue, the system may not have enough available servers to fit this job in service. FCFS stops scheduling additional jobs at this point, even if other jobs in the queue could fit into service. This phenomenon, known as Head-of-the-Line blocking, causes FCFS to underutilize servers, resulting in high mean response time. Although FCFS is a simple policy, analyzing it has been proven difficult due to its dependence on the random arrival order of jobs with different server needs. Only recently, [grosof-2024-marcreset] derived mean response time bounds that confirm the empirical observation that FCFS performs poorly in practice.
@@
Most Servers First (MSF) [bichler, beloglazov, grosof-2023-serverfilling] is a non-preemptive policy that prioritizes jobs with larger server needs. Specifically, whenever the system has available servers, jobs are considered in descending server need order. Jobs that find their required number of servers are put in service successively. To understand both the benefits and the pitfalls of MSF, consider an example where jobs either need one server or $k$ servers. We refer to this case as the one-or-all case for multiserver jobs. In this case, MSF serves jobs in two alternating phases. First, MSF serves $k$-server jobs until none remain in the system. Then MSF serves 1-server jobs until none remain before returning to serve $k$-server jobs. We will show in Section 5.1 that, by switching between these two phases, MSF achieves optimal long-run average resource utilization in the one-or-all case.
@@
While one might hope that MSF leverages its high resource utilization to achieve low mean response time, we also find that MSF takes an increasingly long time to switch phases as the job arrival rate increases (see Section 5.3). This creates a feedback loop in the system whereby many 1-server jobs accumulate while the system processes $k$-server jobs, leading to a long period of serving 1-server jobs during which many $k$-server jobs will accumulate. The two things to note about this process are that (i) despite its name, MSF can spend long periods of time giving priority to 1-server jobs over $k$-server jobs and (ii) because the class of jobs not in service accumulates quickly, there are almost always a large number of jobs in the system under MSF. Figure 1 illustrates this problem via simulations that track the number of jobs in the system under MSF for the one-or-all case. As MSF alternates between phases, jobs of the opposite class accumulate quickly in the queue. While all jobs are eventually served, this behavior ensures that a significant fraction of arriving jobs have long queueing times, leading to a high overall mean response time.
(a) Number of light jobs vs time
(b) Number of heavy jobs vs time
@@
Figure 1: Number of jobs in the system under MSF and MSFQ where there are $k$=32 servers, 90% of the job arrivals are 1-server jobs, mean job sizes are 1 for 1-server jobs and $k$-server jobs, and jobs arrive at a rate of 7.5 jobs/second.
@@
First-Fit is a variant of FCFS that avoids head-of-line blocking by continuing to examine jobs in arrival order after finding a job that does not fit in service. One might hope that this policy gets a near-optimal resource utilization without the harmful periodic behavior of MSF. Unfortunately, in the one-or-all case, First-Fit has the same alternating behavior as MSF, but First-Fit spends even more time serving 1-server jobs.
@@ 1.2 A New Approach: Most Servers First with Quickswap
Our central observation is that, while MSF achieves high resource utilization, it does a poor job on switching which class of jobs are in service. Specifically, in the one-or-all case, MSF ties the decision to switch from serving 1-server jobs to serving $k$-server jobs to the time required to drain all 1-server jobs from the system. This time, which is essentially a partial busy period in an $M/M/k$ queue, explodes as the arrival rate increases or as $k$ becomes large [artalejo2001analysis], allowing many $k$-server jobs to accumulate in the system. To reduce mean response time, our goal is to maintain the high utilization of MSF while shortening the time the scheduling policy takes to switch between job classes.
With this motivation in mind, we propose a new class of policies called Most Servers First with Quickswap (MSFQ), which is designed to improve the performance of MSF in the one-or-all case. Unlike MSF, which tries to drain all 1-server jobs from the system before switching, MSFQ switches to serving $k$-server jobs whenever the number of 1-server jobs in the system falls below a threshold, $\ell$. That is, when the number of 1-server jobs falls below $\ell$, MSFQ stops admitting jobs into service and lets all running 1-server jobs complete. MSFQ can then begin serving $k$-server jobs. By increasing $\ell$, we can shorten the time MSFQ requires to switch phases, reducing the mean response time compared to MSF. Furthermore, we will prove that MSFQ maintains the high resource utilization achieved by MSF.
Figure 1 compares the number of jobs in the system under the MSF and MSFQ policies (the yellow and blue curves, respectively). Setting a sufficiently high value of $\ell$ for MSFQ ($\ell=k-1$ in this case) greatly dampens the feedback that causes jobs to accumulate under MSF. As a result, an MSFQ policy can achieve a much lower mean response time than MSF. We will also show that MSFQ policies are much better at balancing the mean response time between different job classes. While our MSFQ policies are tailored specifically to the one-or-all case, we will also explore generalizations of this policy to an arbitrary number of job classes.
@@ 1.3 Contributions
Inspired by the high resource utilization of the MSF policy, this paper formally defines and analyzes the class of MSFQ policies for scheduling multiserver jobs. Our analysis has two main goals. First, we aim to prove stability results about MSFQ policies. We say the system is stable under a scheduling policy if the policy achieves sufficiently high resource utilization such that the mean number of jobs in the system and the mean response time across jobs are both finite. Second, we will analyze the mean response time under MSFQ policies and their variants both in theory and in simulation to show the advantages of these policies compared to prior approaches. Specifically, the contributions of this paper are as follows.
@@
  • In Section 4.2 we introduce the MSFQ policies, which uses the Quickswap mechanism to force the policies to switch phases faster. Crucially, in Section 5.1 we show that any MSFQ policy matches the resource utilization of MSF by proving that MSFQ is throughput-optimal. Here, throughput-optimality means that MSFQ will stabilize the system whenever the system can be stabilized.
  • Then, in Section 5.3, we analyze the mean response time under MSFQ by approximating the Laplace Transform of the phase durations and number of jobs at the beginning of each phase. While the MSFQ system resembles a polling system (See Section 2.3), MSFQ cannot be analyzed using existing results from the polling literature. Hence, our response time analysis of MSFQ also represents a new contribution to the extensive body of work on polling systems.
  • Finally, Section 6 examines generalizations of MSFQ to real-world settings where jobs’ server needs can vary widely. We evaluate two generalizations of MSFQ, called Static Quickswap and Adaptive Quickswap, and evaluate these policies in simulation using traces from the Google Borg cluster scheduler [tirmazi2020borg]. Our simulations show that a well-designed Quickswap policy can improve mean response time by orders of magnitude. Furthermore, Quickswap policies tend to achieve an equitable mean response time between the job classes as compared to a less fair priority policy like MSF.
@@ 2.1 Systems for Multiserver Job Scheduling
Modern data centers schedule multiserver jobs across thousands of machines, supporting workloads with diverse server needs and job sizes [slurm, tirmazi2020borg, yarn]. None of these schedulers make formal performance guarantees about system stability or mean response time, generally relying on heuristics to make scheduling decisions.
SLURM [slurm] is an open-source cluster scheduler used in data centers and high-performance computing environments. It uses a combination of heuristics and a variant of FCFS scheduling called BackFilling. While this approach can improve resource utilization by running low-priority jobs opportunistically, it requires accurate predictions of job sizes to work well, and can therefore suffer from low resource utilization in practice. Borg [tirmazi2020borg], Google’s internal resource management system for data centers, schedules batch jobs by placing incoming jobs in an FCFS queue. Once there is enough capacity to serve the next batch job, complex heuristics are used to assign the job to a specific set of servers. YARN [yarn], integrated with Hadoop, supports both FCFS and other heuristic policies aimed at optimizing fairness instead of stability or mean response time. Hence, many systems designed to schedule multiserver jobs stand to benefit from improved scheduling policies that are accompanied by formal performance guarantees.
@@ 2.2 Multiserver Job Scheduling in Theory
Prior work from the theory community on multiserver jobs has mostly focused on the stability and response time analysis of FCFS. The stability region of FCFS was studied in [rumyantsev-2017, morozov, afanaseva-2020] in the case where all job sizes follow the same exponential distribution. Subsequently,  [grosof-2023-mama, olliaro-2023] considered the case where jobs belong to one of two job classes, deriving explicit expressions for the stability region of FCFS. For many years, mean response time analysis of FCFS was restricted to systems with just two servers [brill-1984, karatza-2007]. However,  [grosof-2024-marcreset] recently derived explicit bounds on mean response time that are tight up to an additive constant. Matrix geometric approaches [anggraito-mascots-2024, anggraito-cox-2025] have also recently been used to characterize the performance of FCFS systems with two job classes under specific service time distributions. While FCFS is becoming well-understood, all of these analyses confirm that it can perform poorly in terms of both stability and mean response time.
There is comparatively little work on more complex and efficient policies that do not require job preemptions. For example, the well-studied MaxWeight policy is throughput optimal, but requires preemption and is computationally costly to implement in practice [maguluri-2012]. Other recent work on scheduling multiserver jobs has also been restricted to the case of preemptible jobs [wcfs, grosof-2022-pomacs]. There are two prominent examples of throughput-optimal, non-preemptive policies for scheduling multiserver jobs. First, Randomized Timers is a throughput-optimal policy based on MaxWeight that is non-preemptive [psychas2018randomized]. Unfortunately, there is no known mean response time analysis of Randomized Timers, and the policy has been shown to perform poorly in practice. Second, [chen2025improving] recently analyzed a new class of non-preemptive policies called Markovian Service Rate (MSR) policies. An MSR policy precomputes a set of schedules with high resource utilization, and switches between schedules according to a continuous-time Markov chain that is independent of the system state (i.e., queue lengths). The class of MSR policies is throughput-optimal and admits an analysis of mean response time. However, because MSR policies do not consider queue length when switching schedules, they waste capacity unnecessarily, resulting in high mean response time. We will show that MSFQ can significantly outperform MSR policies by considering queue length when switching schedules.